OpenStudy (anonymous):

Find the sum . 7+(-2)

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Note that: \[7+(-2)=7+(-1)\cdot 2=7-2\]

OpenStudy (anonymous):

thanks